OpenStudy (kayders1997):

Please help, gravel is being dumped from a conveyor belt at a rate of 30 ft^3/min, and its coarseness is such that it forms a pile in the shape of a cone whose base diameter and height are always equal. How fast is the height of the pile increasing when the pile is 10 feet high?

OpenStudy (dogzcatz):

Volume of cone = 1/12 x pi x d^2 x h when the height & diameter is 10 =>V = 1/12 x pi x 10 x 10 x 10 = 11000/42 = 261.90 ft^3 In the next minute V = 261.90 + 30 = 291.90 ft^3,let the hight now is h(=d)=>291.90 = 1/12 x 22/7 x h^3 =>h^3 = 1114.55 =>h = 10.37 ft Increase = 0.37 ft/min 6/(5 x 3.14) = 6/15.7 = 0.38

OpenStudy (kayders1997):

1/12?

OpenStudy (dogzcatz):

is it multiple choice?

OpenStudy (kayders1997):

no that is the right answer, but it has to be from calculus standpoint

OpenStudy (dogzcatz):

ooohhh V = (1/3)πr²h h = 2r dV/dt = 30 cubic feet per minute V = (1/3)πh³/4 V = (1/12)πh³ dV/dt = (3/12)πh²(dh/dt) dV/dt = (1/4)πh²(dh/dt) (dV/dt)(4)/πh² = dh/dt 30(4)/π(10)² = dh/dt 120/100π = dh/dt 6/5π = dh/dt

OpenStudy (kayders1997):

Sorry I'm trying to follow this

OpenStudy (kayders1997):

Do r would be 5 correct?

OpenStudy (dogzcatz):

yes

OpenStudy (kayders1997):

And I thought you would have to differentiate h because it's asking for dh/dt so that would be all that's left like h would be gone? Or can't you do that because there asking for the amount at a given time?

OpenStudy (kayders1997):

@dogzcatz

OpenStudy (dogzcatz):

because of given time I can't really do that

OpenStudy (kayders1997):

Oh okay

OpenStudy (kayders1997):

Thanks for help

OpenStudy (kayders1997):

@zepdrix

OpenStudy (kayders1997):

I'm just confused where you got the h^3?

OpenStudy (kayders1997):

@freckles

OpenStudy (kayders1997):

I'm confused

zepdrix (zepdrix):

sup? :O

OpenStudy (kayders1997):

Confused I need an explanation I'm so bad at Related Rate problems :(

OpenStudy (kayders1997):

I think your trying to find dh/dt

zepdrix (zepdrix):

Are you ok with this diagram?

OpenStudy (kayders1997):

Yes

zepdrix (zepdrix):

Volume of a cone is given by,\[\large\rm V=\frac13 \pi r^2h\]In this problem, since the diameter is the height h, the radial length is always half of that,\[\large\rm V=\frac13 \pi \left(\frac{h}{2}\right)^2h\]

OpenStudy (kayders1997):

I plugged in 5 not h/2 that might help

zepdrix (zepdrix):

Maybe we simplify before we take our derivative,\[\large\rm V=\frac{\pi}{12}h^3\]Ok with that step? :o Lemme know what you think.

OpenStudy (kayders1997):

Ughhhhh

zepdrix (zepdrix):

No no no, the height is `changing`. It is varying, a variable. We can't plug anything in for it before differentiating. Our derivative measures change, we can't hold h constant.

OpenStudy (kayders1997):

pi/12? Am I missing something And I'm so bad at that

zepdrix (zepdrix):

\[\large\rm \frac{1}{3}\pi \color{orangered}{\left(\frac{h}{2}\right)^2}h\quad=\quad \frac{\pi}{3}\color{orangered}{\frac{h^2}{2^2}}h\]From that point, since multiplication is commutative, we can sort of move things around how we want. (Keeping denominators as denominators of course).

OpenStudy (kayders1997):

Ohhhhhhhhh duhhhhhh

zepdrix (zepdrix):

We can take a derivative from this point,\[\large\rm V=\frac{\pi}{12}h^3\]Are you ok with doing that step? :) Power rule and chain rule business ya?

OpenStudy (kayders1997):

Yay :)

zepdrix (zepdrix):

I prefer using the prime notation, but whatever works for you,\[\large\rm V'=\frac{3\pi}{12}h^2h'\]And this is the information we started with: \(\rm h=10\) \(\rm V'=30\) Solve for \(\rm h'\)

OpenStudy (kayders1997):

Just a second

OpenStudy (kayders1997):

Okay

zepdrix (zepdrix):

Ya, so plug the stuff in, solve... DO IT! :) Related rates are awesome sauce, you gotta keep studying, they're so great XD

OpenStudy (kayders1997):

6/5pi

zepdrix (zepdrix):

6/(5pi) good good good :)

OpenStudy (kayders1997):

Can we do one more? I really want to do well on my text and youn explain really well

zepdrix (zepdrix):

Open it in another thread please? I'll try to come look at it, but I'm really busy right now :D

OpenStudy (kayders1997):

Sure! It's okay if you don't respond right away, I'll be up a while

zepdrix (zepdrix):

k cool c: